A sphere of radius 1.00 cm is placed in the path of a parallel beam of light of large aperture. The intensity of the light beam is 0.50 W/cm2. If the sphere completely absorbs the radiation falling on it, find the force exerted by the light beam on the sphere.

Grade:12

1 Answers

AJIT AskiitiansExpert-IITD
13 years ago

Dear Ria ,

Change in momentum of photon due to absorbtion  =  h /λ 

no of photons per unit time   =  p/(hc/λ )      p  = power of beam

therefore change in momentum per unit time  =  force  =  h /λ  * p/(hc/λ )      =  p /c 

p = intensity * area =  .5 * 4*pie*r2  =   .5 *4*3.14*(1)2  =  6.28 W

so force  =  6.28 /3 *108 =  2.09 * 10-8 N . The answer of this question in some books as found by me is incorrect.

Please feel free to post as many doubts on our discussion forum as you can. We are all IITians and here to help you in your IIT JEE preparation.


Now you can win exciting gifts by answering the questions on Discussion Forum. So help discuss any query on askiitians forum and become an Elite Expert League askiitian.

Win exciting gifts by answering the questions on Discussion Forum